Can I get help with these please.
4 4/5 / 8/15
5 6/7 + 4 5/8
2 3/8 + 3 3/4
4 5/9 - 7/8
8 2/7 - 3 1/3

Answer by stanbon(57361)

4 4/5 / 8/15 = (24/5)/(8/15) = (24/5)*(15/8) = 3*3 = 9
------------------------------------
5 6/7 + 4 5/8 = (41/7)+(37/8) = (8*41+7*37)/(7*8) = 587/56
------------------------------------
2 3/8 + 3 3/4 = (19/8) + (15/4) = (19 + 30)/8 = 49/8
----------------------------------------
4 5/9 - 7/8 = (41/9) - (7/8) = (8*41-9*7)/(9*8) = 262/75
---------------------
8 2/7 - 3 1/3 = (58/7) - (10/3) = (3*58-7*10)/(7*3) = 104/21
==============
Cheers,
Stan H.